OpenStudy (anonymous):

Any one willing to show me step by step how to draw a parabola with an exquation? (I have one.)

OpenStudy (anonymous):

It depends... what is the equation...

OpenStudy (anonymous):

y=-3/2x+7/2

OpenStudy (anonymous):

Is that supposed to be x^2, if not, how is it a parabola? That would be a straight line :).

OpenStudy (anonymous):

Ohh wait, yeat it is a line. My bad.

OpenStudy (anonymous):

You only need 2 to 3, at most, points to plot a straight line. Thus, you can pick a set of numbers that x is between do you know how to read the inequality: \[-2 \le x \le 2\] And then find your values for y

OpenStudy (anonymous):

 You can draw up a table like the one above to find the y values... or you can do it in your head.